Fermi National Laboratory


Accelerator Update

Wednesday January 21
The day shift began with Operations monitoring store 3183, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 10 AM, the high voltage technicians switched the power for CHL to a different transformer. At 11:21 AM, the techs restored power for most of Meson. Eight Meson buildings will remain on generator power until the feeder repairs finish.

Meson power (feeder 30) faulted at noon. The high voltage techs switched Meson to feeder 31 at 2:14 PM.

A Main Injector (MI) expert began conducting slipstacking studies 2:51 PM. She finished at 5 PM.

At 3:49 PM, feeder 31 faulted off. Meson has no power, except for those buildings with power generators.

The I- Source tripped off at 4:30 PM. It reset okay.

The I- chopper tripped off at 5:05 PM due to a tripped vacuum gauge. It reset okay.

At 7:50 PM, Operations halted stacking so Switchyard experts could conduct an hour of studies.

Stacking resumed at 9:14 PM.

Thursday January 22
The midnight shift began with Operations monitoring store 3183, with Pbar stacking, with feeder 31 still off and generators providing power to certain buildings, and with MiniBooNE taking beam.

At 2:05 AM, Operations terminated store 3183.

Operations put the Tevatron (TeV) into shot setup at 2:49 AM and established store 3185 at 4:58 AM. Its initial luminosity was 42.36E30.

At 6:38 AM, an operator discovered that a MI Lambertson power supply (LAM41) wasn't ramping. After replacing a control card operators had the supply ramping normally by 7:21 AM.

The Collider Detector Facility (CDF) reported seeing a beam problem at 9:31 AM. Operation investigated and discovered that a TeV RF station (TRF2) had sparked. The station remained on.

At 10:24 AM, a Linac expert reported that he had to change out one of the power amplifiers for a Linac RF station (LRF5). He completed the work and had the station back on by 11 AM.

At 12:13 PM, a TeV expert completed moving a sector E3 collimator. This helped clean up the beam for CDF.

MI experts began damper studies at 2:24 PM.

At 2:43 PM, Linac experts switched from the I- Source to the H- Source.

Recycler experts began taking shots of antiprotons from MI around 4:33 PM. They were completed around 5:16 PM.

A Linac RF station (LRF4) tripped off at 4:34 PM for the first of four times over the next two hours.

A Booster power supply (EAPS) tripped off at 6:07 PM due to a water flow fault. Operators called in experts to replace the flow meters. They had the flow meters replaced by 8:53 PM.

At 6:41 PM, Operations noticed a problem with the Pbar Lithium Lens cooling water conductivity. Experts came in to investigate and discovered a bad readback wire at 8:16 PM. They repaired the wire.

Pbar resumed stacking at 8:55 PM.

At 9:47 PM, a Linac RF station (KRF5) began sparking and tripping off. This continued until an expert helped operators fix a regulation problem at 10:37 PM.

At 10:39 PM, at Linac quadrupole power supply (QPS104) tripped off. Operators changed and swapped control cards until 11:46 PM, when it finally stayed on.

The alarm screens at the MCR began showing errors at 11:47 PM.

Friday January 23
The midnight shift began with Operations monitoring store 3185, with Pbar stacking, with MininBooNE taking beam, with the Recycler maintaining a stash of antiprotons, and with operators trying to discover the problem with their alarm screen system.

At 12:50 AM, operator discovered a bad VME crate that was swamping the system with errors. But there were still problems, so Operations called an expert for help. He had the screen system acting normally by 1:37 AM.

Operations noticed at 5:22 AM that the Pbar intensity was dropping. Operators traced the problem back to the H- Source. Their investigations continued.

At 7:14 AM, Operation terminated store 3185.

A Linac expert tuned the H- Source at 7:42 AM and had its intensity up to normal.

The TeV suffered a sector C4 quench at 7:47 AM. The quench was due to a lead flow problem.

Today's Plans
The plan for today is to shoot antiprotons from the Recycler into the Tevatron this morning.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/23/2004   email Fermilab